I need help with this problem, and I also don't really understand multi step equations.
13=5-13+3a

Found 2 solutions by elima, jazzypooh37:
Answer by elima(1433)

13=5-13+3a
First you can subtract 5-13;
You have:
13= -8+3a
Now you want to get the 3a alone; so add 8 to both sides, because what you do to one side you do to the other.
13+8=3a
21 =3a
Now you want to get a by itself; so you will divide 3 on both sides;
21/3=3a/3 Remember what you do to one side you have to do to the other.
7 = a
check:
13=5-13+3a
13=5-13+3(7)
13=-8+21
13=13
You always need to remember the outcome is to find out what the variable equals,
In this case it was a. So you want to do each step until you have the a by itself.

Answer by jazzypooh37(1)